OpenStudy (mitchi35):

Evaluate l 12 l - l -8 l please. :)

Vocaloid (vocaloid):

hint: l -8 l = 8

OpenStudy (mitchi35):

-4????

OpenStudy (mitchi35):

Or 20?

OpenStudy (mitchi35):

Idk what I'm doing

Vocaloid (vocaloid):

| 12 l - l -8 l = 12-8 = ?

OpenStudy (mitchi35):

Either positive or negative 4

OpenStudy (anonymous):

If that is absolute value, the negative and positive does not matter. So for this you take away the negative sign. So what is 12-8?

OpenStudy (mathstudent55):

If you have 12 books and you take away 8 books, do you end up with 4 books or with -4 books?

OpenStudy (mitchi35):

4 @ Kayterbug.cx